OpenStudy (anonymous):

Which equation can be used to model simple harmonic motion? A. d=acos(wt) B. d=asin(wt)+k C. d=cos(awt) D. d=acos[w(t+k)] ****My answer: A is that right?? :)

OpenStudy (anonymous):

sorry... here are the answer choices... :) answer choices A,B,C,D from top to bottom :)

OpenStudy (anonymous):

what do y'all think? is A right? :)

OpenStudy (anonymous):

no i don't mean you, i mean the person that made the question is making me sad

OpenStudy (anonymous):

hahaha ohhh i see.. lol idk... :/ do u get it? does it look like i got it right?? i'm not 100% sure :/

OpenStudy (anonymous):

i debate the wording of this question, but either way http://en.wikipedia.org/wiki/Harmonic_oscillator#Damped_harmonic_oscillator

OpenStudy (espex):

Simple harmonic motion is Asin(omega t)

OpenStudy (anonymous):

i would say d

OpenStudy (anonymous):

how come? :/

OpenStudy (anonymous):

oh we have a conflict.

OpenStudy (anonymous):

hmm? ;/

OpenStudy (anonymous):

@eSpeX you need a phase shift somewhere in there ... .

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (espex):

Well for the available answers, 'b' takes the shift into account.

OpenStudy (anonymous):

that's a vertical shift ...

OpenStudy (anonymous):

not a horizontal shift

OpenStudy (anonymous):

if k in d is φ then its right

OpenStudy (espex):

Then it has to be 'A' since the cos of theta is a shift of 90 from sin.

OpenStudy (anonymous):

sorry without the parenthese

OpenStudy (anonymous):

i'm confused lol :/

OpenStudy (anonymous):

examploes of different shifts http://i.imgur.com/OWWlTo0.png

OpenStudy (anonymous):

d is right if it is d=acos(wt + k) if k is φ

OpenStudy (anonymous):

@eSpeX no wt cannot have a shift in it. w and t are expressions of a variable. variables cannot have operators inside in this

OpenStudy (anonymous):

thats like saying x=y+2 in a taylor series expansion

OpenStudy (anonymous):

Simple Harmonic Motion is A i believe

OpenStudy (espex):

@doppler, but 'd' isn't wt+k, it is w(d+k)

OpenStudy (anonymous):

@AnElephant you're not including a shift

OpenStudy (espex):

@RONNCC A is cos(wt) which is sin(wt + 90)

OpenStudy (anonymous):

yeah only a can be right

OpenStudy (anonymous):

ueh i give up just put A. it depends on the kind of harmonic motion. the simplest simplest case would have the shift as 0 as you're taking it from where you want. thats why this question is bad

OpenStudy (anonymous):

A is the basic form of a harmonic motion curve...

OpenStudy (anonymous):

ooohhh okay.. so the answer is A? :/ and this question is officially and poorly made question? LOL

OpenStudy (anonymous):

*Simple harmonic motion curve

OpenStudy (anonymous):

just say A

OpenStudy (anonymous):

oh okay hahaha thanks for all the help!!! i really appreciate it :)

OpenStudy (anonymous):

technically A and D are the same thing

OpenStudy (anonymous):

but D simplifies to A

OpenStudy (anonymous):

assuming no shift

OpenStudy (anonymous):

quoting textbook ''the complete general solution is x=A1cosωt + A2sinωt''

OpenStudy (anonymous):

ohh they are? weird!!!

OpenStudy (anonymous):

''if you have A1=Acosφ and Α2=-Αsinφ then substitution gives x=Acos(ωt + φ)''

OpenStudy (anonymous):

so it is A with φ=0 but still a little vague as ron said

OpenStudy (anonymous):

@iheartfood cos(x) = sin(x + Pi/2)

OpenStudy (anonymous):

Yeah, my old textbook does not account for phase shifts in the simple harmonic motion chapter

OpenStudy (anonymous):

oh okay.... thanks a bunch y'all :) yup i agree... this is vague :(

OpenStudy (anonymous):

correct mathematically wrong physically